# Irregular Tensor Low-Rank Representation for Hyperspectral Image Representation

### v1

@@ III PROPOSED METHOD
The framework of the proposed method is shown in Fig. 2. First, we use the superpixel segmentation method to divide the origin data into several irregular 3D cubes. Then we complete the irregular 3D data cubes with black shape-complementary data cubes to constitute regular tensor patches. Afterwards, we design a novel low-rank representation model with a non-convex tensor norm to pursue the low-rank representation of the 3D patches that only takes the original irregular 3D data cube into account. To enhance the overall discriminative ability, we introduce a negative nuclear norm term. The details of the proposed method are given as follows.
@@ III-A Low-rank Representation for Irregular Tensor with A Non-convex Tensor Norm
As depicted in Fig. 3, we define $\mathcal{X}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ as an observed hyperspectral image, which usually suffers from spectral variations. The regions that have undergone spectral alterations are indicated by the black dots within the figure. Following [superPCA], we initially apply principal component analysis (PCA) to reduce the original dataset to dimensions of $n_{1}\times n_{2}\times 1$ and adopt entropy rate superpixel segmentation (ERS) [ERS] to segment the compressed PCA data. By utilizing the segmentation results as indices, we can divide the original data into multiple irregular 3D cubes, i.e., $\mathcal{X}=\mathcal{X}_{1}\oplus\mathcal{X}_{2}\oplus\cdots\mathcal{X}_{i}\oplus\cdots\oplus\mathcal{X}_{n}$, where $n$ is the number of superpixels, $\mathcal{X}_{i}$ is denoted for the $i$-th irregular tensor and the symbol $\oplus$ denotes the operation that merges multiple irregular 3D patches into a single regular 3D patch. Given that the shape is irregular, we define the size of each irregular tensor based on the maximum distance between its length ($w$) and width ($b$). Accordingly, the $i$-th irregular tensor is denoted as $\mathcal{X}_{i}\in\mathbb{R}^{w_{i}\times b_{i}\times n_{3}}$.
@@
Although $\mathcal{X}_{i}$ can capture irregular distributions of the input, we cannot directly apply the existing tensor low-rank norms on them to relieve the spectral variation problem. To this end, we introduce a complementary irregular tensor $\mathcal{C}_{i}\in\mathbb{R}^{w_{i}\times b_{i}\times n_{3}}$ to fill in the blanks. Specifically, we shown in Fig. 3, the defined novel operation symbol $\oplus$ can merge two irregular 3D patches into a regular 3D patch. Through this transformation, we obtain a combined regular tensor $\mathcal{X}_{i}\oplus\mathcal{C}_{i}$, where we can perform low-rank approximation on it to extract low-rank representations corresponding to the original irregular tensor $\mathcal{X}_{i}$.
@@
To make the low-rank approximation only affected by the original 3D cubes $\mathcal{X}_{i}$, not by the complementary 3D cube $\mathcal{C}_{i}$, we first define the following notations. Let $\mathcal{L}_{i}^{o}$ and $\mathcal{S}_{i}^{o}$ represent the low-rank and sparse components of $\mathcal{X}_{i}$, respectively, i.e., $\mathcal{X}^{i}=\mathcal{L}_{i}^{o}+\mathcal{S}_{i}^{o}$. Similarly, we define $\mathcal{L}_{i}^{c}$ and $\mathcal{S}_{i}^{c}$ as the low-rank and sparse components of the complementary tensor $\mathcal{C}_{i}$, i.e., $\mathcal{C}^{i}=\mathcal{L}_{i}^{c}+\mathcal{S}_{i}^{c}$. Then, we apply tensor low-rank constraint on the combined regular tensor. i.e., on both $\mathcal{L}_{i}^{o}$ and $\mathcal{L}_{i}^{c}$, while different sparse constraints on $\mathcal{S}_{i}^{o}$ and $\mathcal{S}_{i}^{c}$, i.e., we only impose sparse constraint on $\mathcal{S}_{i}^{o}$, and impose no constraint on $\mathcal{S}_{i}^{c}$. This difference in constraints allows the sparse part of $\mathcal{C}_{i}$ to remain free, while its low-rank part is guided by the low-rank component of $\mathcal{X}_{i}$. As a result, the feature space of the low-rank part of $\mathcal{C}_{i}$ will align with that of the low-rank part of $\mathcal{X}_{i}$, thereby avoiding the influence of the complementary tensor $\mathcal{C}_{i}$ on $\mathcal{X}_{i}$, i.e., only the original irregular plays a role in low-rank tensor representation, while the added complementary irregular tensor does not affect the low-rank property of the original data. Based on this approach, we formulate the following optimization model:
$\min_{[\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c},\mathcal{S}_{i}^{c}]}$ $\|\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}\|_{*}+\lambda_{i}\|\mathcal{S}_{i}^{o}\|_{1}$ (5)
@@
where $\|\cdot\|_{*}$ is a tensor nuclear norm defined in Definition 1, $\lambda_{i}$ is the regularization parameter of the $i$-th superpixel. As suggested by [2020TRPCA], we can set $\lambda_{i}=\alpha/\sqrt{\max(w_{i},b_{i})n_{3}}$ , where $\alpha$ is a variable related to tensor sparsity.
The tensor nuclear norm in Definition 1 regularizes all singular values of tensor data equally, which is not a perfect approximation of the rank function [efficient_of_nonconvex]. To achieve a better approximation of rank function, we introduce a non-convex tensor norm as defined as:
$\|\mathcal{L}\|_{p}=\sum_{i=1}^{n_{3}}\|{\overline{\mathbf{L}}}^{(i)}\|_{p}=\sum_{i=1}^{n_{3}}\sum_{j=1}^{l}\sigma_{j}(\overline{{\mathbf{L}}}^{(i)})^{p}.$ (6)
Here, $p(0<p<1)$ is the power of the singular value and $l=\min(n_{1},n_{2})$. When $p$ is equal to 1, Eq. (6) is equivalent to the tensor nuclear norm defined in Definition 1. For $p$ smaller than 1, this tensor norm will become non-convex but closer to the tensor rank. So the low-rank representation for irregular tensor with a non-convex tensor norm is formulated as follows:
$\min_{[\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c},\mathcal{S}^{o}]}\sum_{i=1}^{n}(\|\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}\|_{p}+\lambda_{i}\|\mathcal{S}_{i}^{o}\|_{1})$ (7)
@@
$\qquad\mathcal{X}=\mathcal{X}_{1}\oplus\mathcal{X}_{2}\oplus\dotsc\oplus\mathcal{X}_{n}$
$\qquad\mathcal{L}^{o}=\mathcal{L}^{o}_{1}\oplus\mathcal{L}^{o}_{2}\oplus\dotsc\oplus\mathcal{L}^{o}_{n}$
$\qquad\mathcal{S}^{o}=\mathcal{S}^{o}_{1}\oplus\mathcal{S}^{o}_{2}\oplus\dotsc\oplus\mathcal{S}^{o}_{n}$
no constraint on $\qquad\text{no constraint on }\mathcal{S}^{c}_{i},\forall i.$
@@ III-B Promoting the Discriminative Ability
However, the above model processes the irregular 3D data cubes individually, overlooking that a superpixel always contains more than one category of materials as shown in Fig. 2 (II). When seeking low-rank representation in the local area, the non-dominant class in one superpixel block will be over penalized to make it close to the dominant one, which leads to a decrease in discriminability between different categories in one superpixel and decrease in similarity among pixels belonging to same category in different superpixels. For example, in Fig. 2 (II), we can see that blocks $B$, $E_{1}$ and $E_{2}$ are the non-dominant classes within each superpixel and blocks $E_{1}$ and $E_{2}$ originate from one category in the original data. But after executing Eq. (7), they are over penalized, thereby pushed closer to the categories that occupy the dominant position in irregular 3D data cubes as shown in Fig. 2 (IV). This will decrease the discriminability among different categories in one superpixel (e.g., $A$ and $B$, $E_{1}$ and $C$, $E_{2}$ and $D$). And the similarity of the pixels belonging to one category also decreases (e.g., $E_{1}$ and $E_{2}$).
To address this issue, we introduce a global regularization term $-\|\mathcal{L}^{o}\|_{\star}:=-\|\mathbf{L}_{(3)}\|_{*}$, where $\mathbf{L}_{(3)}\in\mathbb{R}^{n_{1}n_{2}\times n_{3}}$ is the third order unfolding $\mathcal{L}^{o}$ and $\|\cdot\|_{*}$ is the nuclear norm of the 2D matrix. Different from the low-rank prior, the global term aims to promote the discriminability among different categories and improve the consistency of the same category in the global area by increasing the singular values of the whole matrix. The final objective function of ITLRR is formulated as:
$\min_{[\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c},\mathcal{S}^{o}]}\sum_{i=1}^{n}\left(\|\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i}\|_{p}+\lambda_{i}\|\mathcal{S}^{o}_{i}\|_{1}\right)-\beta\|\mathcal{L}^{o}\|_{\star}$ (8)
[…]
$\qquad\mathcal{X}=\mathcal{X}_{1}\oplus\mathcal{X}_{2}\oplus\dotsc\oplus\mathcal{X}_{n}$
$\qquad\mathcal{L}^{o}=\mathcal{L}^{o}_{1}\oplus\mathcal{L}^{o}_{2}\oplus\dotsc\oplus\mathcal{L}^{o}_{n}$
$\qquad\mathcal{S}^{o}=\mathcal{S}^{o}_{1}\oplus\mathcal{S}^{o}_{2}\oplus\dotsc\oplus\mathcal{S}^{o}_{n}$
no constraint on $\qquad\text{no constraint on }\mathcal{S}^{c}_{i},\forall i,$
where $\beta\geq 0$ serves as a trade-off parameter. By balancing the global term and the local irregular low-rank term, the dissimilarity among pixels belonging to different categories are expected to increase, while that of the pixels belonging to the same category will decrease.
@@ III-C Optimization
$\min_{[\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c},\mathcal{S}^{o}]}$ $\sum_{i=1}^{n}\left(\|\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i}\|_{p}+\lambda_{i}\|\mathcal{S}^{o}_{i}\|_{1}\right)-\beta\|\mathcal{L}^{o}\|_{\star}$ (9)
@@
where $\mathcal{Y}=\mathcal{Y}_{1}\oplus\mathcal{Y}_{2}\oplus\dotsc\oplus\mathcal{Y}_{n}$, $\mathcal{Y}_{i}\in\mathbb{R}^{w_{i}\times b_{i}\times n_{3}}$ serves as the Lagrangian multiplier for the $i$-th superpixel, and $\mu>0$ acts as the penalty parameter. Note that there are no constraints on $\mathcal{C}_{i}$ and $\mathcal{S}^{c}_{i}$, and we only need to solve for $\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}$ and $\mathcal{S}^{o}$. We solve it by an alternative manner, i.e., alternatively update one variable and fix others.
1) The $\mathcal{L}$ sub-problem can be expressed as:
$\underset{[\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}]}{\operatorname{min}}$ $\sum_{i=1}^{n}\|\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i}\|_{p}-\beta\|\mathcal{L}^{o}\|_{\star}$ (10)
where the second term is concave, making it challenging to solve. In order to simplify Eq. (10), we first linearize the concave term $-\|\mathcal{L}^{o}\|_{\star}$ through its first-order Taylor expansion around $\mathcal{L}^{o^{t}}$, which was obtained in the preceding iteration, with $t\geq 0$ denoting the iteration index. This approximation is expressed as follows:
@@
where $\mathcal{T}^{t}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ represents the sub-gradient of $\|\mathcal{L}^{o}\|_{\star}$ evaluated at $\mathcal{L}^{o^{t}}$, defined as $\mathcal{T}^{t}=\partial\|\mathcal{L}^{o^{t}}\|_{\star}:=\mathscr{F}_{3}(\partial\|\mathbf{L}_{(3)}^{o^{t}}\|_{*})$. The explicit formulation of $\partial\|\cdot\|_{*}$ is detailed in Theorem 2.
@@
$\min_{[\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}]}g(\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i})$ $=\sum_{i=1}^{n}\|\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i}\|_{p}$ (12)
$\quad-\beta\|\mathcal{L}^{o^{t}}\|_{\star}-\beta\langle\mathcal{T}^{o^{t}},\mathcal{L}^{o}-\mathcal{L}^{o^{t}}\rangle$
$\quad+\frac{\mu}{2}\left\|\mathcal{L}^{o}-\left(\mathcal{X}-\mathcal{S}^{o}+\frac{\mathcal{Y}}{\mu}\right)\right\|_{F}^{2}$
$=\sum_{i=1}^{n}\|\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i}\|_{p}+\frac{\mu}{2}\|\mathcal{L}^{o}-\mathcal{A}\|_{F}^{2}$
$\quad-\beta\|\mathcal{L}^{o^{t}}\|_{\star}-\beta\langle\mathcal{T}^{o^{t}},-\mathcal{L}^{o^{t}}\rangle,$
@@
Since $\left\{\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}\right\}_{i=1}^{n}$ are independent, it can be divided into $n$ independent sub-problems. The sub-problem for the $i$-th superpixel is equivalent to:
$\min_{[\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}]}g(\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c})$ $=\|\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i}\|_{p}+\frac{\mu}{2}\|\mathcal{L}_{i}^{o}-\mathcal{A}_{i}\|_{F}^{2}$ (13)
$=\|\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i}\|_{p}$
$\quad+\frac{\mu}{2}\|\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i}-(\mathcal{A}_{i}\oplus\mathcal{L}^{c}_{i})\|_{F}^{2},$
where the introduced $\mathcal{L}^{c}_{i}$ will not affect the Frobenius norm term.
Eq. (13) is equivalent to:
$\min_{\mathcal{N}_{i}}\|\mathcal{N}_{i}\|_{p}+\frac{\mu}{2}\|\mathcal{N}_{i}-\mathcal{M}_{i}\|_{F}^{2},$ (14)
@@
$\min_{\mathcal{N}_{i}}g(\mathcal{N}_{i})$ $=\sum_{j=1}^{n_{3}}\|\overline{\mathbf{N}}_{i}^{(j)}\|_{p}+\frac{\mu}{2}\|\overline{\mathcal{N}_{i}}-\overline{\mathcal{M}_{i}}\|_{F}^{2}$ (15)
$=\sum_{j=1}^{n_{3}}\left(\|\overline{\mathbf{N}}_{i}^{(j)}\|_{p}+\frac{\mu}{2}\|\overline{\mathbf{N}}_{i}^{(j)}-\overline{\mathbf{M}}_{i}^{(j)}\|_{F}^{2}\right).$
@@
$\min_{\overline{\mathbf{N}}_{i}^{(j)}}\|\overline{\mathbf{N}}_{i}^{(j)}\|_{p}+\frac{\mu}{2}\|\overline{\mathbf{N}}_{i}^{(j)}-\overline{\mathbf{M}}_{i}^{(j)}\|_{F}^{2}.$ (16)
According to Theorem 3, the optimal solution to the $j$-th sub-problem of Eq. (16) is
@@
2) The $\mathcal{S}^{o}$ sub-problem can be expressed as follows:
$\sum_{i=1}^{n}\left(\underset{\mathcal{S}_{i}}{\operatorname{min}}\lambda_{i}\|\mathcal{S}^{o}_{i}\|_{1}+\frac{\mu}{2}\left\|\mathcal{S}_{i}-\mathcal{B}_{i}\right\|_{F}^{2}\right),$ (18)
@@
Eq. (18) is a set of $\ell_{1}$ norm minimization problem, which can be solved by the soft-thresholding operator, i.e.,
@@
Initialize $\rho=1.1$, $\mu=1e-10$, $\mu_{\text{max}}=1e10$ and $\epsilon=1e-3$.
@@
Create $\mathcal{L}_{i}^{o}$, $\mathcal{L}_{i}^{c}$, $\mathcal{S}_{i}^{o}$ and $\mathcal{Y}_{i}$ with dimensions equivalent to $\mathcal{X}_{i}$.

### v3

@@ III Proposed Method
While tensor-based methods can effectively preserve the spatial structure of HSIs, they can only be applied to regular 3D data cubes, overlooking the fact that the same material in an HSI is usually distributed within irregular homogeneous regions. To solve this problem, we propose a novel tensor low-rank representation model that can process irregular data cubes and accordingly promote the representation ability.
Fig. 2 illustrates the framework of the proposed method. First, we use a typical superpixel segmentation method to divide the origin data into several irregular 3D cubes. Then we complete the irregular 3D data cubes with black shape-complementary data cubes to constitute regular tensor patches. Afterward, we design a novel low-rank representation model with a non-convex tensor norm to pursue the low-rank representation of the 3D patches that only takes the original irregular 3D data cube into account. To enhance the overall discriminative ability, we introduce a global negative nuclear norm term. The details of the proposed method are given below.
@@ III-A Low-rank Representation for Irregular Tensor with A Non-convex Tensor Norm
As depicted in Fig. 4, we define $\mathcal{X}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ as the input HSI, which usually suffers from spectral variations. The regions that have undergone spectral alterations are indicated by the black dots within the figure. Following [52], we initially apply principal component analysis (PCA) to reduce the original dataset to dimensions of $n_{1}\times n_{2}\times 3$ and adopt entropy rate superpixel segmentation (ERS) [53] to segment the compressed PCA data. By utilizing the segmentation results as indices, we can divide the original data into multiple non-overlapping irregular 3D cubes, i.e., $\mathcal{X}=\mathcal{X}_{1}\oplus\mathcal{X}_{2}\oplus\cdots\mathcal{X}_{i}\oplus\cdots\oplus\mathcal{X}_{n}$, where $n$ is the number of superpixels and $\mathcal{X}_{i}$ is denoted for the $i$-th irregular tensor.
@@
Although $\mathcal{X}_{i}|_{i=1}^{n}$ can capture irregular distributions of the input, we cannot directly apply the existing tensor low-rank norms on them to relieve the spectral variation problem. To this end, we introduce a complementary irregular tensor $\mathcal{C}_{i}$ to fill in the blanks. Specifically, we determine the smallest regular tensor with dimensions $\mathbb{R}^{w_{i}\times b_{i}\times n_{3}}$ that can exactly accommodate $\mathcal{X}_{i}$. The complementary irregular tensor $\mathcal{C}_{i}$ is the region of the regular tensor that is not part of the original irregular tensor. Through introducing $\mathcal{C}_{i}$, we obtain a combined regular tensor $\mathcal{X}_{i}\oplus\mathcal{C}_{i}\in\mathbb{R}^{w_{i}\times b_{i}\times n_{3}}$, where we can perform low-rank approximation on it to extract low-rank representations corresponding to the original irregular tensor.
@@
To make the low-rank approximation only affected by the original 3D cubes $\mathcal{X}_{i}$, not by the complementary 3D cube $\mathcal{C}_{i}$, we define the following notations. Let $\mathcal{L}_{i}^{o}$ and $\mathcal{S}_{i}^{o}$ represent the low-rank and sparse components of $\mathcal{X}_{i}$, respectively, i.e., $\mathcal{X}^{i}=\mathcal{L}_{i}^{o}+\mathcal{S}_{i}^{o}$. Similarly, we define $\mathcal{L}_{i}^{c}$ and $\mathcal{S}_{i}^{c}$ as the low-rank and sparse components of the complementary tensor $\mathcal{C}_{i}$, i.e., $\mathcal{C}^{i}=\mathcal{L}_{i}^{c}+\mathcal{S}_{i}^{c}$. Then, we apply tensor low-rank constraint on the combined regular tensor. i.e., on both $\mathcal{L}_{i}^{o}$ and $\mathcal{L}_{i}^{c}$ ($\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}$), while different sparse constraints on $\mathcal{S}_{i}^{o}$ and $\mathcal{S}_{i}^{c}$, i.e., we only impose sparse constraint on $\mathcal{S}_{i}^{o}$, and impose no constraint on $\mathcal{S}_{i}^{c}$. This difference in constraints allows the sparse part of $\mathcal{C}_{i}$ to remain free, while its low-rank part is guided by the low-rank component of $\mathcal{X}_{i}$. As a result, the feature space of the low-rank part of $\mathcal{C}_{i}$ will align with that of the low-rank part of $\mathcal{X}_{i}$, thereby avoiding the influence of the complementary tensor $\mathcal{C}_{i}$ on $\mathcal{X}_{i}$, i.e., only the original irregular plays a role in low-rank tensor representation, while the added complementary irregular tensor does not affect the low-rank property of the original data. Based on this approach, we formulate the following optimization model for the $i$-th irregular data cube:
$\min_{\{\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c},\mathcal{S}_{i}^{o}\}}$ $\|\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}\|_{*}+\lambda_{i}\|\mathcal{S}_{i}^{o}\|_{1}$ (6)
@@
where $\|\cdot\|_{*}$ is a tensor nuclear norm defined in Definition 1, $\lambda_{i}$ is the regularization parameter of the $i$-th superpixel. As suggested by [18], we can set $\lambda_{i}=\alpha/\sqrt{\max(w_{i},b_{i})n_{3}}$ , where $\alpha$ is a variable related to tensor sparsity.
The tensor nuclear norm in Definition 1 regularizes all singular values of tensor data equally, which is not a perfect approximation of the rank function [26]. To achieve a better approximation of rank function, we introduce a non-convex tensor norm [54] for a tensor $\mathcal{A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ as defined as:
$\|\mathcal{A}\|_{S_{p}}^{p}=\sum_{i=1}^{n_{3}}\|{\overline{\mathbf{A}}}^{(i)}\|_{S_{p}}^{p}=\sum_{i=1}^{n_{3}}\sum_{j=1}^{l}\sigma_{j}(\overline{{\mathbf{A}}}^{(i)})^{p}.$ (7)
Here, $p(0<p\leq 1)$ is the power of the singular value and $l=\min(n_{1},n_{2})$. $\overline{\mathbf{A}}^{(i)}\in\mathbb{R}^{n_{1}\times n_{2}}$ is the $i$-th frontal slice of $\overline{\mathcal{A}}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$, where $\overline{\mathcal{A}}=fft(\mathcal{A},[],3)$. When $p$ is equal to 1, Eq. (7) is equivalent to the tensor nuclear norm defined in Definition 1. For $p$ smaller than 1, this tensor norm will become nonconvex but closer to the tensor rank. So the low-rank representation for irregular tensor with a non-convex tensor norm for the input tensor $\mathcal{X}$ is formulated as follows:
$\min_{\left\{\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c},\mathcal{S}_{i}^{o}\right\}_{i=1}^{n}}\sum_{i=1}^{n}(\|\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}\|_{S_{p}}^{p}+\lambda_{i}\|\mathcal{S}_{i}^{o}\|_{1})$ (8)
@@
no constraint on $\qquad\qquad\text{no constraint on }\mathcal{S}^{c}_{i},\forall i.$
where $\mathcal{X}=\mathcal{X}_{1}\oplus\mathcal{X}_{2}\oplus\dotsc\oplus\mathcal{X}_{n}$, $\mathcal{L}^{o}=\mathcal{L}^{o}_{1}\oplus\mathcal{L}^{o}_{2}\oplus\dotsc\oplus\mathcal{L}^{o}_{n}$ and $\mathcal{S}^{o}=\mathcal{S}^{o}_{1}\oplus\mathcal{S}^{o}_{2}\oplus\dotsc\oplus\mathcal{S}^{o}_{n}$.
@@ III-B Promoting the Discriminative Ability
However, the above model processes the irregular 3D data cubes individually, overlooking that a superpixel potentially contains more than one category of materials as shown in Fig. 2 (II). When seeking low-rank representation in the local area, the non-dominant class in one superpixel block will be over-penalized to make it close to the dominant one, which leads to a decrease in discriminability between different categories in one superpixel and a decrease in similarity among pixels belonging to the same category in different superpixels. For example, in Fig. 2 (II), we can see that blocks $B$, $E_{1}$ and $E_{2}$ are the non-dominant classes within each superpixel and blocks $E_{1}$ and $E_{2}$ originate from one category in the original data. But after executing Eq. (8), they are over-penalized, thereby pushed closer to the categories that occupy the dominant position in irregular 3D data cubes as shown in Fig. 2 (IV). This will decrease the discriminability among different categories in one superpixel (e.g., $A$ and $B$, $E_{1}$ and $C$, $E_{2}$ and $D$). And the similarity of the pixels belonging to one category also decreases (e.g., $E_{1}$ and $E_{2}$).
To address this issue, we introduce a global regularization term $-\|\mathcal{L}^{o}\|_{\star}:=-\|\mathbf{L}_{(3)}\|_{*}$, where $\mathbf{L}_{(3)}\in\mathbb{R}^{n_{1}n_{2}\times n_{3}}$ is the third-order unfolding $\mathcal{L}^{o}$ and $\|\cdot\|_{*}$ is the nuclear norm of the 2D matrix. Different from the irregular tensor low-rank prior, the global term aims to promote discriminability among different categories and improve the consistency of the same category in different superpixels by increasing the singular values of the whole matrix. The final objective function of ITLRR is formulated as:
$\min_{\left\{\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c},\mathcal{S}_{i}^{o}\right\}_{i=1}^{n}}$ $\sum_{i=1}^{n}\left(\|\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}\|_{S_{p}}^{p}+\lambda_{i}\|\mathcal{S}_{i}^{o}\|_{1}\right)-\beta\|\mathcal{L}^{o}\|_{\star}$
s.t. $\mathcal{X}=\mathcal{L}^{o}+\mathcal{S}^{o},\mathcal{C}_{i}=\mathcal{L}_{i}^{c}+\mathcal{S}_{i}^{c},\forall i$ (9)
no constraint on $\text{no constraint on }\mathcal{S}_{i}^{c},\forall i$
where $\beta\geq 0$ serves as a trade-off parameter, $\mathcal{X}=\mathcal{X}_{1}\oplus\mathcal{X}_{2}\oplus\dotsc\oplus\mathcal{X}_{n}$, $\mathcal{L}^{o}=\mathcal{L}^{o}_{1}\oplus\mathcal{L}^{o}_{2}\oplus\dotsc\oplus\mathcal{L}^{o}_{n}$ and $\mathcal{S}^{o}=\mathcal{S}^{o}_{1}\oplus\mathcal{S}^{o}_{2}\oplus\dotsc\oplus\mathcal{S}^{o}_{n}$. By balancing the global term and the local irregular low-rank term, the dissimilarity among pixels belonging to different categories is expected to increase, while that of the pixels belonging to the same category will decrease.
@@ III-C Optimization
[…]
where $\mathcal{Y}=\mathcal{Y}_{1}\oplus\mathcal{Y}_{2}\oplus\dotsc\oplus\mathcal{Y}_{n}$, $\mathcal{Y}_{i}\in\mathbb{R}^{w_{i}\times b_{i}\times n_{3}}$ serves as the Lagrangian multiplier for the $i$-th superpixel, and $\mu>0$ acts as the penalty parameter. Note that there are no constraints on $\mathcal{C}_{i}$ and $\mathcal{S}^{c}_{i}$, and we only need to solve for $\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}$ and $\mathcal{S}^{o}$. We solve Eq. (10) in an alternative manner, i.e., alternatively update one variable and fix others.
1) The $\left\{\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}\right\}_{i=1}^{n}$ sub-problem can be expressed as:
$\underset{\left\{\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}\right\}_{i=1}^{n}}{\operatorname{min}}$ $\sum_{i=1}^{n}\|\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i}\|_{S_{p}}^{p}-\beta\|\mathcal{L}^{o}\|_{\star}$ (11)
@@
where the second term is concave, making it challenging to solve. In order to simplify Eq. (11), we first linearize the concave term $-\|\mathcal{L}^{o}\|_{\star}$ through its first-order Taylor expansion around $\mathcal{L}^{o^{t}}$, which was obtained in the preceding iteration, with $t\geq 0$ denoting the iteration index. The approximation is expressed as follows:
@@
where $\mathcal{T}^{t}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ represents the sub-gradient of $\|\mathcal{L}^{o}\|_{\star}$ [55] evaluated at $\mathcal{L}^{o^{t}}$, defined as $\mathcal{T}^{t}=\partial\|\mathcal{L}^{o^{t}}\|_{\star}:=\mathscr{F}_{3}(\partial\|\mathbf{L}_{(3)}^{o^{t}}\|_{*})$. The explicit formulation of $\partial\|\cdot\|_{*}$ is detailed in Theorem 2.
@@
[…]
@@
Since $\left\{\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}\right\}_{i=1}^{n}$ are independent, Eq. (13) can be divided into $n$ independent sub-problems. The sub-problem for the $i$-th superpixel formulated as:
$\min_{\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c}}g(\mathcal{L}_{i}^{o}\oplus\mathcal{L}_{i}^{c})$ $=\|\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i}\|_{S_{p}}^{p}+\frac{\mu}{2}\|\mathcal{L}_{i}^{o}-\mathcal{A}_{i}\|_{F}^{2}$ (14)
$=\|\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i}\|_{S_{p}}^{p}$
$\quad+\frac{\mu}{2}\|\mathcal{L}^{o}_{i}\oplus\mathcal{L}^{c}_{i}-(\mathcal{A}_{i}\oplus\mathcal{L}^{c}_{i})\|_{F}^{2}.$
Eq. (14) can be simplified to:
$\min_{\mathcal{N}_{i}}\|\mathcal{N}_{i}\|_{S_{p}}^{p}+\frac{\mu}{2}\|\mathcal{N}_{i}-\mathcal{M}_{i}\|_{F}^{2},$ (15)
@@
$\min_{\mathcal{N}_{i}}g(\mathcal{N}_{i})$ $=\sum_{j=1}^{n_{3}}\|\overline{\mathbf{N}}_{i}^{(j)}\|_{S_{p}}^{p}+\frac{\mu}{2}\|\overline{\mathcal{N}_{i}}-\overline{\mathcal{M}_{i}}\|_{F}^{2}$ (16)
$=\sum_{j=1}^{n_{3}}\left(\|\overline{\mathbf{N}}_{i}^{(j)}\|_{S_{p}}^{p}+\frac{\mu}{2}\|\overline{\mathbf{N}}_{i}^{(j)}-\overline{\mathbf{M}}_{i}^{(j)}\|_{F}^{2}\right).$
@@
$\min_{\overline{\mathbf{N}}_{i}^{(j)}}\|\overline{\mathbf{N}}_{i}^{(j)}\|_{S_{p}}^{p}+\frac{\mu}{2}\|\overline{\mathbf{N}}_{i}^{(j)}-\overline{\mathbf{M}}_{i}^{(j)}\|_{F}^{2}.$ (17)
According to Theorem 3, the optimal solution to Eq. (17) is given by:
@@
2) The $\left\{\mathcal{S}_{i}^{o}\right\}_{i=1}^{n}$ sub-problem can be expressed as follows:
$\sum_{i=1}^{n}\left(\underset{\mathcal{S}_{i}^{o}}{\operatorname{min}}\lambda_{i}\|\mathcal{S}^{o}_{i}\|_{1}+\frac{\mu}{2}\left\|\mathcal{S}^{o}_{i}-\mathcal{B}_{i}\right\|_{F}^{2}\right),$ (19)
@@
Eq. (19) is a set of $\ell_{1}$ norm minimization problems, which can be solved by the soft-thresholding operator, i.e.,
@@
Initialize $\rho=1.1$, $\mu=1e-10$, $\mu_{\text{max}}=1e10$ and
$\epsilon=1e-3$.
@@
Create $\mathcal{L}_{i}^{o}$, $\mathcal{L}_{i}^{c}$, $\mathcal{S}_{i}^{o}$ and $\mathcal{Y}_{i}$ with dimensions equivalent
to $\mathcal{X}_{i}$.